# xRFM: Accurate, scalable, and interpretable feature learning models for tabular data

### v1

@@ 5 Features learned by xRFM on tabular data
We use two approaches for identifying features selected by the AGOP. The first is to identify the elements along the diagonal of the AGOP with highest magnitude. By definition, these entries indicate how much a leaf RFM’s predictions vary when perturbing a given coordinate. As such, they are a natural measure of feature importance. The second approach is to examine the loadings onto the top eigenvectors of the AGOP. This approach allows us to identify joint effects of feature perturbations on the prediction. Namely, if coordinate $i$ of the top eigenvector is positive and coordinate $j$ is negative, then increasing one of these features and decreasing the other leads to a change in prediction.
We used both approaches above to understand what features were learned by xRFM on tabular data from TALENT and meta-test benchmarks (Fig. 7). As examples, we study the AGOPs for four such datasets:
  1. (1)
    California housing: a regression task for predicting the average price of a house.
  2. (2)
    Covertype: a multiclass classification task for identifying the dominant tree species in a given location.
  3. (3)
    NYC Taxi Tipping: a regression task for predicting the dollar tip amount in a taxi ride.
  4. (4)
    Breast cancer: a classification task for identifying malignancy from features of biopsy images.
As the California housing and breast cancer datasets contain fewer than 5050k samples (the parameter we used for leaf size), we visualize a single AGOP. For covertype, we visualize one of the AGOPs from a leaf RFM (for this dataset, the leaf RFM AGOPs generally indicated the same pattern), and for taxi tipping we visualize several leaf RFM AGOPs. Upon visualization, it is apparent that AGOPs indicate low rank structure: either they highlight a subset of coordinates relevant for prediction (Fig. 7A, B) or they have a decay in the eigenvalue spectrum (Fig. 7C).
In Fig. 7A and B, we list feature names for the features with highest diagonal AGOP entries (darker shades of red indicate higher values). In all cases, we find that AGOP identified sensible features for prediction. For the California housing dataset, xRFM identified longitude, or east-west location, as the most important feature for predicting the average price of a house. Given that beach fronts (and major cities) in California are typically located to the west, the importance of house longitude is consistent with the hypothesis that homes closer to the beach are more expensive on average. For the Covertype prediction dataset, the example AGOP from a leaf RFM model showed that elevation, distance to roadways, and distance to firepoints were the most important features. This finding is consistent with the hypothesis that different elevations have different climates which significantly affect viability of tree species, and that fires / roadways can select for tree types that are fire-adapted or have avoided human logging. For the taxi tipping dataset, we observed that leaf RFMs identified varying local features. For example, one leaf RFM AGOP (denoted Taxi Tipping 1 in Fig. 7B) selected pickup location as an important feature, while this feature was less important for a different leaf RFM AGOP (Taxi Tipping 3). Furthermore, fare code and the MTA tax have varying feature relationships at each leaf: in leaf RFM 1 and 2, the fare code and MTA tax had neutral or synergistic effects on the prediction value, while for leaf RFM 3, increasing fare code had the opposite effect on prediction as increasing MTA tax (shown as a blue square in Fig. 7B).
For the breast cancer dataset, the most important feature found by xRFM was the average number of concave points in the biopsy image, which has been shown to be a significant indicator of malignancy [25]. When examining the top eigenvector of the AGOP, we find that standard error in the compactness measurements of cell nuclei is the feature with highest importance that has an opposite effect on the malignancy label as concavity (Fig. 7C). This finding suggests that benign cells have less uniform compactness than malignant cells.
@@ Acknowledgments
We would like to thank Jingang Qu, Ingo Steinwart, Tizian Wenzel, and Gaël Varoquaux for relevant discussions. We acknowledge support from the National Science Foundation (NSF) and the Simons Foundation for the Collaboration on the Theoretical Foundations of Deep Learning through awards DMS-2031883 and #814639 as well as the TILOS institute (NSF CCF-2112665), the Office of Naval Research (ONR N000142412631) and DARPA (HR001125CD020). This work used the programs (1) XSEDE (Extreme science and engineering discovery environment) which is supported by NSF grant numbers ACI-1548562, and (2) ACCESS (Advanced cyberinfrastructure coordination ecosystem: services & support) which is supported by NSF grants numbers #2138259, #2138286, #2138307, #2137603, and #2138296. Specifically, we used the resources from SDSC Expanse GPU compute nodes, and NCSA Delta system, via allocations TG-CIS220009.
@@ A.1 Additional leaf-RFM modifications
    We tune over a more general class of kernels $K_{p,q}(x,x^{\prime})=\exp(-\|x-x^{\prime}\|_{q}^{p}/L^{p})$ that are positive definite for $0<p\leq q\leq 2$ [30, Theorems 1, 5]. We typically search over $p\in\mathcal{U}(0.7,1.4)$ and $q\in\{p,2\}$ (Table A.1). In the case $q=p$, we obtain a kernel that is a product of 1D kernels and is therefore not rotation-invariant.
@@
    We tune over whether or not to use adaptive bandwidth, which involves separately scaling $L$ at each leaf RFM by either $\left(\text{Median}(\|x-x^{\prime}\|_{1}))\right)^{-1}$ (if using product kernels) or $\left(\text{Median}(\|x-x^{\prime}\|_{2}))\right)^{-1}$ (if using orthogonally invariant kernels) for $x\neq x^{\prime}$ at every leaf independently. Adaptive scaling allows xRFM to adapt to the variance of covariates at different leaves.
@@ A.3 Hyperparameters
Hyperparameter TALENT Meta-test
Bandwidth $\log\mathcal{U}(1,200)$ $\log\mathcal{U}(0.4,80)$
Bandwidth Mode {constant} {constant, adaptive}
Categorical Transformations {one_hot} {ordinal_encoding, one_hot}
Diagonal {False, True} {False, True}
Early Stop Multiplier 1.06 1.05
Exponent $p$ $\mathcal{U}(0.7,1.4)$ $\mathcal{U}(0.7,1.3)$
Kernel Type $K_{p,p}$, $K_{p,2}$} $K_{p,p}$, $K_{p,2}$}
Normalization {standard} {standard}
Regularization $\log\mathcal{U}(10^{-6},1)$ $\log\mathcal{U}(10^{-5},50)$
Refill size ($N_{\mathrm{val}}$) 1500 1500
Table A.1: Search spaces for xRFM on the TALENT (Figure 3) and Meta-test benchmarks (Figure 6).
@@
For scaling kernel ridge regression and the standard kernel RFM to large TALENT datasets, we used Eigenpro-2 (EP2) [23] that was initialized with the coefficients obtained from directly solving RFM on a random sample of 70,000 points. To tune the optimization hyperparameters for EP2, we tuned the model parameters for 100 trials on a random 70,000 sample subset, then for the final run, we used the tuned hyperparameters (and for kernel RFM the AGOP from the best iteration). For TabPFN-v2, we used the code provided from the benchmark github (https://github.com/LAMDA-Tabular/TALENT), which subsampled to 10,000 samples to avoid out-of-memory issues. The provided code did not apply TabPFN-v2 to datasets with more than 10 classes.
@@ A.4 Algorithmic details
For compute gradients of the predictor on training data (for AGOP computation), we omit the contribution to the gradient from the kernel evaluation between each training point and itself. This is because the kernel function is often not differentiable (for e.g. Laplace kernels) when evaluated for two identical points.
@@
   $X_{M}\leftarrow X\odot\text{diag}(M)^{1/2}_{t}$
@@
   $f^{(t)}(x)=K(x\odot\text{diag}(M)^{1/2}_{t},X_{M})\alpha_{t}$ $\triangleright$ $\odot$ denotes element-wise multiplication
@@
  • $\mathcal{A}$ : linear solver used inside LeafRFM
@@
   $(\alpha_{\ell},M_{\ell})\leftarrow$ LeafRFM($\mathcal{D}_{\ell}^{\mathrm{train}},\mathcal{D}_{\ell}^{\mathrm{val}}$, $K$, $\text{LeafHyp},\mathcal{A}$) $\triangleright$ Alg. A.1
@@ Appendix B All results
xRFM 5.50 0.315 0.041 15.0 41.0 59.0 78.0 0.315
TabPFN-v2 6.46 0.323 0.066 23.0 45.0 61.0 72.0 0.323
CatBoost 7.67 0.336 0.052 9.00 25.0 43.0 69.0 0.336
RealMLP 8.14 0.329 0.075 7.00 22.0 38.0 62.0 0.329
ModernNCA 9.36 0.365 0.096 14.0 30.0 38.0 55.0 0.365
LightGBM 9.82 0.353 0.061 5.00 20.0 35.0 53.0 0.353
XGBoost 10.6 0.355 0.070 2.00 15.0 33.0 47.0 0.355
TabR 10.6 0.365 0.132 5.00 22.0 34.0 47.0 0.365
FTT 12.0 0.388 0.148 3.00 11.0 16.0 34.0 0.388
Laplace KRR 12.3 0.373 0.101 3.00 10.0 20.0 34.0 0.373
MLP-PLR 12.7 0.381 0.141 0.0 9.00 16.0 34.0 0.381
Excelformer 12.7 0.399 0.156 0.0 2.00 11.0 28.0 0.399
PTARL 13.1 0.390 0.113 3.00 6.00 8.00 19.0 0.390
RandomForest 13.3 0.389 0.087 5.00 14.0 22.0 35.0 0.389
AutoInt 14.3 0.399 0.157 1.00 2.00 7.00 13.0 0.399
Node 14.4 0.451 0.195 1.00 7.00 18.0 31.0 0.451
MLP 15.2 0.418 0.167 1.00 3.00 5.00 11.0 0.418
DCN2 15.2 0.473 0.227 0.0 2.00 9.00 20.0 0.473
ResNet 15.6 0.426 0.175 0.0 3.00 7.00 15.0 0.426
Tangos 16.3 0.432 0.173 0.0 2.00 4.00 10.0 0.432
SNN 17.1 0.419 0.174 0.0 2.00 4.00 8.00 0.419
kNN 19.1 0.459 0.189 2.00 4.00 5.00 11.0 0.459
TabNet 21.5 0.475 0.226 0.0 0.0 0.0 0.0 0.475
GrowNet 23.0 0.619 0.323 0.0 0.0 0.0 1.00 0.619
SVM 23.2 0.612 0.345 0.0 0.0 1.00 1.00 0.612
@@
TabPFN-v2 7.04 0.823 0.935 26.5 51.5 61.8 70.6 0.108
xRFM 7.63 0.823 0.921 8.82 38.2 55.9 69.1 0.109
RealMLP 7.70 0.823 0.918 10.3 20.6 45.6 66.2 0.107
TabR 8.07 0.828 0.932 8.82 25.0 38.2 58.8 0.106
ModernNCA 8.90 0.825 0.913 10.3 29.4 42.6 60.3 0.106
CatBoost 10.3 0.819 0.906 1.47 19.1 38.2 51.5 0.117
LightGBM 11.7 0.813 0.883 5.88 20.6 30.9 44.1 0.125
XGBoost 12.2 0.815 0.887 4.41 17.6 26.5 41.2 0.124
ResNet 12.2 0.807 0.879 0.0 4.41 19.1 35.3 0.127
FTT 12.8 0.810 0.868 0.0 4.41 10.3 27.9 0.125
@@
Laplace KRR 13.5 0.804 0.865 5.88 16.2 20.6 33.8 0.133
DCN2 14.0 0.806 0.865 1.47 5.88 10.3 27.9 0.127
MLP 14.2 0.805 0.864 0.0 5.88 8.82 19.1 0.131
SNN 14.9 0.805 0.858 0.0 0.0 2.94 8.82 0.129
AutoInt 15.7 0.804 0.858 0.0 0.0 1.47 8.82 0.130
RandomForest 15.8 0.797 0.832 5.88 8.82 11.8 26.5 0.137
Excelformer 16.1 0.805 0.854 0.0 0.0 5.88 17.6 0.127
Tangos 16.5 0.797 0.844 0.0 1.47 7.35 13.2 0.133
@@
Danets 17.7 0.795 0.829 0.0 1.47 2.94 7.35 0.137
PTARL 19.0 0.796 0.808 0.0 0.0 1.47 4.41 0.139
kNN 20.1 0.786 0.773 4.41 7.35 10.3 14.7 0.152
LogReg 20.1 0.767 0.755 1.47 4.41 7.35 13.2 0.162
TabTransformer 21.5 0.771 0.739 0.0 1.47 1.47 4.41 0.155
@@
GrowNet 24.4 0.691 0.596 0.0 0.0 1.47 5.88 0.216
SwitchTab 25.0 0.748 0.674 0.0 1.47 1.47 1.47 0.187
TabNet 26.0 0.753 0.616 0.0 0.0 0.0 0.0 0.183
NaiveBayes 29.4 0.609 0.328 0.0 0.0 0.0 0.0 0.321
NCM 30.5 0.627 0.279 0.0 0.0 1.47 1.47 0.326
@@
RealMLP 3.67 0.730 0.962 8.33 58.3 91.7 91.7 0.164
xRFM 6.08 0.716 0.941 25.0 25.0 66.7 75.0 0.185
TabR 6.08 0.720 0.945 8.33 50.0 50.0 66.7 0.172
ResNet 6.33 0.720 0.946 0.0 16.7 50.0 83.3 0.178
ModernNCA 6.75 0.713 0.911 33.3 41.7 50.0 83.3 0.168
FTT 8.79 0.709 0.917 0.0 8.33 25.0 50.0 0.187
MLP-PLR 8.79 0.712 0.923 0.0 0.0 25.0 50.0 0.185
MLP 9.58 0.710 0.919 0.0 0.0 8.33 41.7 0.190
Laplace KRR 10.8 0.679 0.887 0.0 16.7 41.7 50.0 0.218
DCN2 11.2 0.707 0.913 0.0 8.33 8.33 8.33 0.193
AutoInt 12.1 0.701 0.897 0.0 0.0 0.0 25.0 0.194
SNN 12.2 0.704 0.904 0.0 0.0 0.0 33.3 0.197
CatBoost 12.8 0.679 0.840 8.33 33.3 33.3 33.3 0.218
Danets 15.9 0.687 0.880 0.0 0.0 0.0 0.0 0.215
Tangos 16.0 0.660 0.820 8.33 8.33 8.33 16.7 0.225
XGBoost 16.2 0.678 0.872 8.33 8.33 8.33 25.0 0.236
Excelformer 16.5 0.675 0.849 0.0 0.0 0.0 16.7 0.217
TabCaps 16.8 0.670 0.851 0.0 0.0 0.0 0.0 0.231
LightGBM 16.9 0.668 0.848 0.0 0.0 0.0 16.7 0.261
PTARL 18.6 0.656 0.821 0.0 0.0 0.0 0.0 0.233
kNN 20.2 0.637 0.799 0.0 0.0 0.0 8.33 0.273
RandomForest 20.7 0.620 0.780 0.0 8.33 8.33 8.33 0.311
TabTransformer 20.8 0.582 0.700 0.0 0.0 8.33 8.33 0.307
LogReg 22.0 0.538 0.620 0.0 0.0 0.0 0.0 0.338
SVM 25.0 0.486 0.522 0.0 8.33 8.33 8.33 0.385
SwitchTab 25.2 0.525 0.602 0.0 0.0 0.0 0.0 0.372

### v2

@@ 5 Features learned by xRFM on tabular data
We use two approaches for identifying features selected by the AGOP. The first is to identify the elements along the diagonal of the AGOP with highest magnitude. By definition, these entries indicate how much a leaf RFM’s predictions vary when perturbing a given coordinate. As such, they are a natural measure of feature importance. The second approach is to examine the loadings onto the top eigenvectors of the AGOP. This approach allows us to identify joint effects of feature perturbations on the prediction. Namely, if coordinate $i$ of the top eigenvector is positive and coordinate $j$ is negative, then increasing one of these features and decreasing the other changes the prediction.
We use both approaches above to understand what features are learned by xRFM on tabular data from scikit-learn and meta-test (Fig. 7). As examples, we study the AGOPs for four such datasets: (1) California housing - a regression task for predicting the average price of a house, (2) Covertype - a multiclass classification task for identifying the dominant tree species in a given location, (3) NYC Taxi Tipping - a regression task for predicting the dollar tip amount in a taxi ride, and (4) Breast cancer - a classification task for identifying malignancy from features of biopsy images.
As the California housing and breast cancer datasets contain fewer than 5050k samples (the parameter we used for leaf size), we visualize a single AGOP. For covertype, we visualize one of the AGOPs from a leaf RFM (for this dataset, the leaf RFM AGOPs generally indicate the same pattern), and for taxi tipping we visualize several leaf RFM AGOPs. Upon visualization, it is apparent that AGOPs indicate low rank structure: either they highlight a subset of coordinates relevant for prediction (Fig. 7A, B) or they have a decay in the eigenvalue spectrum (Fig. 7C).
In Fig. 7A and B, we list feature names for the features with highest diagonal AGOP entries (darker shades of red indicate higher values). In all cases, we find that AGOP identifies sensible features for prediction. For the California housing dataset, xRFM identifies longitude, or east-west location, as the most important feature for predicting the average price of a house. Given that beach fronts (and major cities) in California are typically located to the west, the importance of house longitude is consistent with the hypothesis that homes closer to the beach are more expensive on average. For the Covertype prediction dataset, the example AGOP from a leaf RFM shows that elevation, distance to roadways, and distance to firepoints are the most important features. This finding is consistent with the hypothesis that elevations with different climates and the existence of fires / roadways can significantly affect viability of different tree species. For the taxi tipping dataset, we observe that leaf RFMs identify varying local features. For example, one leaf RFM (denoted Taxi Tipping 1 in Fig. 7B) selected pickup location as an important feature, while this feature was less important for a different leaf RFM (Taxi Tipping 3). Furthermore, fare code and the MTA tax have varying feature relationships at each leaf: in leaf RFM 1 and 2, the fare code and MTA tax has neutral or synergistic effects on the prediction value, while for leaf RFM 3, increasing fare code has the opposite effect on prediction as increasing MTA tax (shown as a blue square in Fig. 7B).
@@
For the breast cancer dataset, the most important feature found by xRFM is the mean number of concave points in the biopsy image, which has been shown to be a significant indicator of malignancy [27]. When examining the top eigenvector of the AGOP, we find that standard error in the compactness measurements of cell nuclei is the feature with highest importance and has an opposite effect on malignancy as concavity (Fig. 7C). Our method suggests that benign cells have less uniform compactness than malignant cells.
@@ Acknowledgments
We would like to thank Jingang Qu, Ingo Steinwart, Tizian Wenzel, and Gaël Varoquaux for relevant discussions. We thank Charles Durham for contributing the KerMac library, which led to drastic speed up in kernel evaluations and gradient computations. We acknowledge support from the National Science Foundation (NSF) and the Simons Foundation for the Collaboration on the Theoretical Foundations of Deep Learning through awards DMS-2031883 and #814639 as well as the TILOS institute (NSF CCF-2112665), the Office of Naval Research (ONR N000142412631) and DARPA (HR001125CD020). This work used the programs (1) XSEDE (Extreme science and engineering discovery environment) which is supported by NSF grant numbers ACI-1548562, and (2) ACCESS (Advanced cyberinfrastructure coordination ecosystem: services & support) which is supported by NSF grants numbers #2138259, #2138286, #2138307, #2137603, and #2138296. Specifically, we used the resources from SDSC Expanse GPU compute nodes, and NCSA Delta system, via allocations TG-CIS220009.
@@ A.1 Additional leaf-RFM modifications
    We tune over a more general class of kernels $K_{p,q}(x,x^{\prime})=\exp(-\|x-x^{\prime}\|_{p}^{q}/L^{q})$ that are positive definite for $0<q\leq p\leq 2$ [32, Theorems 1, 5]. We typically search over $p\in\mathcal{U}(0.7,1.4)$ and $p\in\{q,2\}$ (Table A.1).
@@
    We tune over whether or not to use adaptive bandwidth, which involves separately scaling $L$ at each leaf RFM by $\left(\text{Median}(\|x-x^{\prime}\|_{p}))\right)^{-1}$ for $x\neq x^{\prime}$ at every leaf independently. Adaptive scaling allows xRFM to adapt to the variance of covariates at different leaves.
@@ A.3 Hyperparameters
Hyperparameter TALENT TabArena-Lite Meta-test
Bandwidth $\log\mathcal{U}(1,200)$ $\log\mathcal{U}(0.5,200)$ $\log\mathcal{U}(0.4,80)$
Bandwidth Mode {constant} {constant} {constant, adaptive}
Categorical Transformations {one_hot} {one_hot} {ordinal_encoding, one_hot}
Diagonal {False, True} {False, True} {False, True}
Early Stop Multiplier 1.1 1.1 1.1
Exponent $q$ $\mathcal{U}(0.7,1.4)$ $\mathcal{U}(0.7,1.4)$ $\mathcal{U}(0.7,1.3)$
[…]
$p$ (when kernel type is $K_{p,q}$) $\mathcal{U}(q,q+0.8(2-q))$ $\mathcal{U}(q,q+0.8(2-q))$ $\mathcal{U}(q,q+0.8(2-q))$
Normalization {standard} {standard} {standard}
Regularization $\log\mathcal{U}(10^{-6},1)$ $\log\mathcal{U}(10^{-6},10)$ $\log\mathcal{U}(10^{-5},50)$
Refill size ($N_{\mathrm{val}}$) 1500 1500 1500
Min. split subset size ($C$) 60,000 60,000 60,000
Table A.1: Search spaces for xRFM on the TALENT (Figure 3), TabArena-Lite (Figure 3), and Meta-test benchmarks (Figure 6).
@@
For scaling kernel ridge regression and the standard kernel RFM to large TALENT datasets, we used Eigenpro-2 (EP2) [25] that was initialized with the coefficients obtained from directly solving RFM on a random sample of 70,000 points. To tune the optimization hyperparameters for EP2, we tuned the model parameters for 100 trials on a random 70,000 sample subset, then for the final run, we used the tuned hyperparameters (and for kernel RFM the AGOP from the best iteration). For TabPFN-v2, we used the code provided from the benchmark github (https://github.com/LAMDA-Tabular/TALENT), which subsampled to 10,000 samples to avoid out-of-memory issues. The provided code did not apply TabPFN-v2 to datasets with more than 10 classes. For efficient $L_{p}^{q}$ kernel computations on GPU we used the KerMac library (https://github.com/Kernel-Machines/kermac).
@@ A.4 Algorithmic details
For computing gradients of the predictor on training data (for AGOP computation), we omit the contribution to the gradient from the kernel evaluation between each training point and itself. This is because the kernel function is often not differentiable (e.g. Laplace kernels) when evaluated for two identical points.
@@
   $X_{M}\leftarrow X\odot\text{diag}(M_{t})^{1/2}$
@@
   $f^{(t)}(x)=K(x\odot\text{diag}(M_{t})^{1/2},X_{M})\alpha_{t}$ $\triangleright$ $\odot$ denotes element-wise multiplication
@@
   $(\alpha_{\ell},M_{\ell})\leftarrow$ LeafRFM($\mathcal{D}_{\ell}^{\mathrm{train}},\mathcal{D}_{\ell}^{\mathrm{val}}$, $K$, LeafHyp) $\triangleright$ Alg. A.1
@@ Appendix B All results
xRFM 4.70 0.311 0.036 20.0 56.0 69.0 84.0 0.311
TabPFN-v2 6.55 0.323 0.067 20.0 45.0 57.0 72.0 0.323
CatBoost 7.79 0.336 0.053 7.00 24.0 41.0 68.0 0.336
RealMLP 8.20 0.329 0.076 8.00 21.0 37.0 60.0 0.329
ModernNCA 9.39 0.365 0.097 14.0 30.0 38.0 57.0 0.365
LightGBM 9.91 0.353 0.062 6.00 16.0 34.0 52.0 0.353
XGBoost 10.6 0.355 0.071 2.00 12.0 32.0 47.0 0.355
TabR 10.7 0.365 0.133 5.00 20.0 34.0 46.0 0.365
FTT 12.0 0.388 0.149 3.00 11.0 17.0 33.0 0.388
Laplace KRR 12.3 0.373 0.102 3.00 11.0 21.0 35.0 0.373
MLP-PLR 12.7 0.381 0.142 0.0 8.00 15.0 34.0 0.381
Excelformer 12.7 0.399 0.157 0.0 2.00 9.00 28.0 0.399
PTARL 13.1 0.390 0.114 2.00 6.00 9.00 20.0 0.390
RandomForest 13.3 0.389 0.088 5.00 12.0 22.0 35.0 0.389
AutoInt 14.3 0.399 0.158 1.00 2.00 6.00 12.0 0.399
Node 14.4 0.451 0.195 0.0 6.00 18.0 31.0 0.451
MLP 15.2 0.418 0.167 1.00 3.00 4.00 11.0 0.418
DCN2 15.3 0.473 0.227 0.0 2.00 10.0 20.0 0.473
ResNet 15.6 0.426 0.175 0.0 4.00 8.00 15.0 0.426
Tangos 16.3 0.432 0.173 0.0 1.00 4.00 9.00 0.432
SNN 17.2 0.419 0.174 0.0 2.00 4.00 8.00 0.419
kNN 19.1 0.459 0.190 2.00 4.00 4.00 11.0 0.459
TabNet 21.5 0.475 0.227 0.0 0.0 0.0 0.0 0.475
GrowNet 23.0 0.619 0.323 0.0 0.0 0.0 0.0 0.619
SVM 23.2 0.612 0.346 0.0 0.0 1.00 1.00 0.612
@@
TabPFN-v2 7.15 0.823 0.935 25.0 51.5 63.2 70.6 0.108
xRFM 7.60 0.825 0.933 11.8 29.4 48.5 66.2 0.107
RealMLP 7.64 0.823 0.918 10.3 20.6 48.5 66.2 0.107
TabR 7.98 0.828 0.932 8.82 27.9 41.2 60.3 0.106
ModernNCA 8.88 0.825 0.912 10.3 30.9 44.1 61.8 0.106
CatBoost 10.3 0.819 0.905 1.47 19.1 36.8 51.5 0.117
LightGBM 11.7 0.813 0.883 5.88 22.1 33.8 44.1 0.125
XGBoost 12.1 0.815 0.886 4.41 19.1 26.5 41.2 0.124
ResNet 12.3 0.807 0.879 0.0 4.41 16.2 36.8 0.127
FTT 12.8 0.810 0.868 0.0 5.88 10.3 29.4 0.125
@@
Laplace KRR 13.5 0.804 0.864 5.88 16.2 20.6 33.8 0.133
DCN2 14.0 0.806 0.865 1.47 4.41 10.3 26.5 0.127
MLP 14.2 0.805 0.864 0.0 4.41 8.82 19.1 0.131
SNN 14.9 0.805 0.858 0.0 0.0 4.41 8.82 0.129
AutoInt 15.7 0.804 0.858 0.0 0.0 1.47 7.35 0.130
RandomForest 15.7 0.797 0.831 5.88 8.82 13.2 27.9 0.137
Excelformer 16.1 0.805 0.853 0.0 0.0 4.41 17.6 0.127
Tangos 16.5 0.797 0.844 0.0 1.47 4.41 13.2 0.133
@@
Danets 17.7 0.795 0.829 0.0 1.47 2.94 5.88 0.137
PTARL 19.1 0.796 0.808 0.0 1.47 1.47 4.41 0.139
kNN 20.1 0.786 0.772 2.94 7.35 10.3 14.7 0.152
LogReg 20.2 0.767 0.755 1.47 4.41 7.35 13.2 0.162
TabTransformer 21.5 0.771 0.739 0.0 1.47 1.47 2.94 0.155
@@
GrowNet 24.4 0.691 0.596 0.0 0.0 2.94 5.88 0.216
SwitchTab 25.0 0.748 0.674 0.0 1.47 1.47 2.94 0.187
TabNet 26.0 0.753 0.615 0.0 0.0 0.0 0.0 0.183
NaiveBayes 29.4 0.609 0.327 0.0 0.0 0.0 0.0 0.321
NCM 30.5 0.627 0.278 0.0 0.0 1.47 1.47 0.326
@@
RealMLP 3.75 0.730 0.964 8.33 50.0 91.7 91.7 0.164
xRFM 5.46 0.718 0.949 25.0 50.0 66.7 83.3 0.183
TabR 6.00 0.720 0.948 16.7 50.0 50.0 66.7 0.172
ResNet 6.33 0.720 0.949 0.0 16.7 50.0 83.3 0.178
ModernNCA 6.83 0.713 0.913 33.3 41.7 50.0 83.3 0.168
FTT 8.79 0.709 0.919 0.0 8.33 25.0 50.0 0.187
MLP-PLR 8.79 0.712 0.925 0.0 0.0 25.0 50.0 0.185
MLP 9.58 0.710 0.922 0.0 0.0 8.33 41.7 0.190
Laplace KRR 10.8 0.679 0.889 0.0 16.7 41.7 50.0 0.218
DCN2 11.2 0.707 0.915 0.0 8.33 8.33 8.33 0.193
AutoInt 12.2 0.701 0.899 0.0 0.0 0.0 16.7 0.194
SNN 12.2 0.704 0.906 0.0 0.0 0.0 33.3 0.197
CatBoost 12.8 0.679 0.842 8.33 25.0 33.3 33.3 0.218
Danets 15.9 0.687 0.882 0.0 0.0 0.0 0.0 0.215
Tangos 16.1 0.660 0.823 0.0 8.33 8.33 16.7 0.225
XGBoost 16.2 0.678 0.874 8.33 8.33 8.33 25.0 0.236
Excelformer 16.6 0.675 0.851 0.0 0.0 0.0 16.7 0.217
TabCaps 16.8 0.670 0.853 0.0 0.0 0.0 0.0 0.231
LightGBM 16.9 0.668 0.850 0.0 0.0 0.0 16.7 0.261
PTARL 18.7 0.656 0.823 0.0 0.0 0.0 0.0 0.233
kNN 20.2 0.637 0.800 0.0 0.0 0.0 8.33 0.273
RandomForest 20.7 0.620 0.782 0.0 8.33 8.33 8.33 0.311
TabTransformer 20.8 0.582 0.702 0.0 0.0 8.33 8.33 0.307
LogReg 22.0 0.538 0.622 0.0 0.0 0.0 0.0 0.338
SVM 25.1 0.486 0.523 0.0 8.33 8.33 8.33 0.385
SwitchTab 25.2 0.525 0.603 0.0 0.0 0.0 0.0 0.372